Problem 3.00E5-Calculating the velocity of a moving object A ball is thrown downward with an initial velocity of 15 m/s. Using the approximate value of g-10 m/s2 calculate the velocity of the The velocity of the ball ism/s.
Answer #1

Solution)

We know,

V = u+at

here u = initial velocity = 15 m/s

Given, g=10 m/s^2 = acceleration due to gravity

t = 3.2 sec

final velocity V = 15 +(10 * 3.2) = 47 m/s (Ans)

=======